# supplement

### v1

@@ 4.1 Univariate functional time-series (UFTS) forecasting method
where $\mu_{s}^{g}(u)=\text{E}\{\mathcal{X}_{t,s}^{g}(u)\}$. Via Mercer’s lemma, there exists a set of orthonormal sequence $(\phi_{k})$ of continuous functions in $\mathcal{L}^{2}(\mathcal{I})$ and a non-increasing sequence $\lambda_{k,s}^{g}$ of positive numbers, such that
@@
$=\mu_{s}^{g}(u)+\sum_{k=1}^{K}\beta_{t,k,s}^{g}\phi_{k,s}^{g}(u)+e_{t,s}^{g}(u),$
where the principal component score $\beta_{t,k,s}^{g}$ is given by the projection of $[\mathcal{X}_{t,s}^{g}(u)-\mu_{s}^{g}(u)]$ in the direction of the $k$th eigenfunction $\phi_{k,s}^{g}(u)$, i.e., $\beta_{t,k,s}^{g}=\langle\mathcal{X}_{t,s}^{g}(u)-\mu_{s}^{g}(u),\phi_{k,s}^{g}(u)\rangle$; and $e_{t,s}^{g}(u)$ denotes the model error function with a mean of zero and a finite variance, and $K<n$ is the number of retained components.
@@
$K=\operatorname*{arg\,min}_{k=1,\dots,k_{\max}}\left\{\frac{\widehat{\lambda}_{k+1,s}^{g}}{\widehat{\lambda}_{k,s}^{g}}\times\mathds{1}\Big{\{}\frac{\widehat{\lambda}_{k,s}^{g}}{\widehat{\lambda}_{1,s}^{g}}\geq\eta\Big{\}}+\mathds{1}\Big{\{}\frac{\widehat{\lambda}_{k,s}^{g}}{\widehat{\lambda}_{1,s}^{g}}<\eta\Big{\}}\right\},$ (3)
where $\mathds{1}\{\cdot\}$ denotes the binary indicator function, $\eta=\frac{1}{\ln[\max(\widehat{\lambda}_{1,s}^{g},n)]}$ is a small positive value. Instead of searching through $n$, we restrict our searching range by setting $k_{\max}=\#\{k|\widehat{\lambda}_{k,s}^{g}\geq\frac{1}{n}\sum^{n}_{k=1}\widehat{\lambda}_{k,s}^{g},k\geq 1\}$. For comparison, we also consider $K=6$ as suggested in Hyndman et al. (2013).
@@
$\widehat{\mathcal{X}}_{n+h|n,s}^{g}(u)$ $=\text{E}\big{[}\mathcal{X}_{n+h,s}^{g}(u)|\bm{\mathcal{X}}_{s}^{g}(u),\bm{B}\big{]}$
$=\widehat{\mu}_{s}^{g}(u)+\sum^{K}_{k=1}\widehat{\beta}_{n+h|n,k,s}^{g}\widehat{\phi}_{k,s}^{g}(u),$
@@ 4.2 Multivariate functional time-series (MFTS) forecasting method
The univariate functional time-series method does not incorporate a correlation between female and male data within the same region. However, explicitly modeling the correlation between multiple series may improve forecast accuracy. We consider a multivariate functional time-series method to jointly model and forecast multiple series that could be correlated with each other.
Let $\mathcal{X}_{t,s}^{F}(u)$ and $\mathcal{X}_{t,s}^{M}(u)$ represent female and male subnational age distribution of death counts. As our multiple functional time series have the same function support, we consider data where each observation consists of $\omega\geq 2$ functions, that is, $[\mathcal{X}_{t,s}^{F}(u),\mathcal{X}_{t,s}^{M}(u)]\in\mathds{R}^{2}$, where $u\in\mathcal{I}$.
The multivariate functional time series are stacked in a vector. Let $\nu^{F}_{s}(u)$ and $\nu^{M}_{s}(u)$ represent the mean function. For $u,v\in\mathcal{I}$, the theoretical cross-covariance function can be defined with elements
@@
where $\bm{\mathcal{X}}_{s}(u)=[\bm{\mathcal{X}}_{s}^{F}(u),\bm{\mathcal{X}}_{s}^{M}(u)]$, $\bm{\mathcal{X}}_{s}^{F}(u)=[\mathcal{X}_{1,s}^{F}(u),\dots,\mathcal{X}_{n,s}^{F}(u)]$ denote stacked historical functions, and $\bm{\nu}_{s}(u)=[\bm{\nu}_{s}^{F}(u),\bm{\nu}_{s}^{M}(u)]$. Moreover, $\bm{\beta}_{s}=(\bm{\beta}_{1,s},\dots,\bm{\beta}_{K,s})$ and $\bm{\beta}_{1,s}=(\beta_{1,1,s},\dots,\beta_{n,1,s})$ is the vector of principal component scores, and
$\bm{\Phi}_{s}(u)=\left[\begin{array}[]{cccccc}\phi_{1,s}^{F}(u)&\cdots&\phi_{K,s}^{F}(u)&&\cdots&\\ &\cdots&&\phi_{1,s}^{M}(u)&\cdots&\phi_{K,s}^{M}(u)\end{array}\right]_{2\times(K\times 2)}$
@@
Conditioning on the past functions $\bm{\mathcal{X}}_{s}(u)$ and the estimated functional principal components $\bm{\Phi}_{s}(u)$, the $h$-step-ahead point forecast of $\bm{\mathcal{X}}_{n+h,s}(u)$ can be expressed as
@@ 4.3 Multilevel functional time-series (MLFTS) forecasting method
The multilevel functional data model strongly resembles to the two-way functional analysis of variance studied by Morris & Carroll (2006) and Cuesta-Albertos & Febrero-Bande (2010). The basic idea is to extract common patterns shared by multiple series $R_{t,s}(u)$ and series-specific patterns $U_{t,s}^{g}(u)$. The common trend and series-specific trend are modeled by projecting them onto the eigenvectors of covariance functions of the aggregated and series-specific stochastic processes, relatively. For $t=1,\dots,n$, a curve can be expressed as
@@
The sample versions of series-specific mean function, a common trend and series-specific residual trend for a set of functional time series can be estimated by
@@
Plugging (5) to (7) into (4), we obtain
@@
When the common trend can explain the primary mode of total variability, the value of the within-cluster variability is close to 1.
@@
4.4 Two-way functional analysis of variance
To model $\mathcal{X}_{t,s}^{g}(u)$, we consider a two-way functional ANOVA decomposition, expressed as:
@@ 4.4 Two-way functional analysis of variance
The functional ANOVA model can be estimated by means with
@@
Some identifiability constraints exist, $\forall u\in\mathcal{I}$, $\sum^{n_{s}}_{s=1}\alpha_{s}(u)=\sum^{2}_{g=1}\beta^{g}(u)=0$, and $\sum^{n_{s}}_{s=1}\mathcal{X}_{t,s}^{g}(u)=\sum^{2}_{g=1}\mathcal{X}_{t,s}^{g}(u)=0$, $\forall t$.
The residual functions $\widehat{\epsilon}_{t,s}^{g}(u)=\mathcal{X}_{t,s}^{g}(u)-\widehat{\mu}(u)-\widehat{\alpha}_{s}(u)-\widehat{\beta}^{g}(u)$ are time-varying, for $t=1,\dots,n$. Following Jiménez-Varón et al. (2024), we model and forecast the residual functions via the MFTS.
@@ 4.5 Two-stage functional principal component analyses
    The first functional principal component scores from each of $P$ sets of functional time series are combined into an $P\times 1$ vector. We fit factor models to the scores to further reduce the dimension into an $r\times 1$ vector, where $r<<P$. The same is carried out for the second, third, and so on until the $p_{0}$th scores. The vector of $P$ functional time series is reduced to an $r\times p_{0}$ matrix consisting of factors.
@@
    A univariate time series model can be fitted to each factor and forecasts are generated. The forecast factors can be used to construct forecast functions.
@@ 5 Construction of prediction intervals
We equally split the data sample consisting of 48 years from 1975 to 2022 into training, validation, and testing sets, each consisting of 16 years. Using the data in the training sample, we implement an expanding window forecast scheme to obtain the $h$-step-ahead density forecasts in the validation set for $h=1,2,\dots,15$. For each horizon, we have different numbers of curves in the validation set. For instance, when $h=1$, we have 16 years to evaluate the forecast errors, denoted by $\widehat{\epsilon}_{m,s}^{g}(u)=d_{m,s}^{g}(u)-\widehat{d}_{m,s}^{g}(u)$ for $m=1,2,\dots,M=16$; when $h=15$, we have two years to evaluate the residual functions between the samples in the validation set and their corresponding forecasts. From these residual functions, we compute the functional standard deviation, denoted by
@@
$-\underline{\xi}_{\alpha}\gamma(u)\leq\widehat{\epsilon}_{m,s}^{g}(u)\leq\overline{\xi}_{\alpha}\gamma(u),$
@@
Typically, the constants $\overline{\xi}_{\alpha}$ and $\underline{\xi}_{\alpha}$ are chosen equal. By a law of large numbers, one may achieve
@@
where $\mathds{1}(\cdot)$ denotes binary indicator function and $M$ denotes the number of years in the validation set.
@@
In Appendix C, we also present the conformal prediction interval approach of Shafer & Vovk (2008) that does not require the tuning parameter. Essentially, by considering a quantile of the absolute residual functions, the pointwise prediction interval for the out-of-sample testing set is centered around the point forecast, with its spread determined by the chosen quantile.
@@ 6.1 Expanding window scheme
An expanding window analysis of a time-series model is commonly used to assess model and parameter stability and prediction accuracy over time. The expanding window analysis determines the constancy of a model’s parameter by computing parameter estimates and their forecasts over an expanding window through the sample (for details, see Zivot & Wang, 2006, pp. 313–314).
Using the first 32 years from 1975 to 2006 in the Japanese subnational age- and sex-specific life-table death counts, we produce one- to 16-step-ahead forecasts. Through an expanding-window approach, we re-estimate the parameters in the time-series forecasting models using the first 33 years from 1975 to 2007. Forecasts from the estimated models are then produced for one- to 15-step-ahead forecasts. We iterate this process by increasing the sample size by one year until reaching the end of the data period in 2022. This process produces 16 one-step-ahead forecasts, 15 two-step-ahead forecasts, $\dots$, and one 16-step-ahead forecast. We compare these forecasts with the holdout samples to determine the out-of-sample forecast accuracy. In Figure 4, we display a diagram of the expanding window forecast scheme for forecast horizon $h=1$, although we also consider other forecast horizons from $h=2$ to $16$.
TimeTrainFTrainFTrainFTrainF1975:20061975:2007$\vdots$1975:2021Training WindowForecast (F)
@@ 6.2 Point forecast error metrics
Since life-table death counts can be considered a probability density function, we utilize some density evaluation measures. These measures include the discrete version of the KLD (Kullback & Leibler, 1951), the square root of the JSD (Shannon, 1948). For two probability density functions, denoted by $d_{m+\xi}$ and $\widehat{d}_{m+\xi}$, the symmetric version of the KLD, also known as Jeffrey divergence, is defined as
$\text{KLD}(h)=$ $\text{D}_{\text{KL}}(d_{m+\xi}||\widehat{d}_{m+\xi})+\text{D}_{\text{KL}}(\widehat{d}_{m+\xi}||d_{m+\xi})$
=\ = $\frac{1}{111\times(17-h)}\sum^{16}_{\xi=h}\sum_{x=1}^{111}d_{m+\xi}(u_{x})\cdot\left[\ln d_{m+\xi}(u_{x})-\ln\widehat{d}_{m+\xi}(u_{x})\right]+$
$1$×($\frac{1}{111\times(17-h)}\sum^{16}_{\xi=h}\sum_{x=1}^{111}\widehat{d}_{m+\xi}(u_{x})\cdot\left[\ln\widehat{d}_{m+\xi}(u_{x})-\ln d_{m+\xi}(u_{x})\right],$ (9)

### v2

@@ 4.1 Univariate functional time-series (UFTS) forecasting method
where $\mu_{s}^{g}(u)=\text{E}\{\mathcal{X}_{t,s}^{g}(u)\}$. By Mercer’s lemma, there exists a set of orthonormal sequences $(\phi_{k})$ of continuous functions in $\mathcal{L}^{2}(\mathcal{I})$ and a non-increasing sequence $\lambda_{k,s}^{g}$ of positive numbers such that
@@
$=\mu_{s}^{g}(u)+\sum_{k=1}^{K_{s}^{g}}\beta_{t,k,s}^{g}\phi_{k,s}^{g}(u)+e_{t,s}^{g}(u),$
where the principal component score $\beta_{t,k,s}^{g}$ is given by the projection of $\mathcal{X}_{t,s}^{g}(u)-\mu_{s}^{g}(u)]$ in the direction of the $k$th eigenfunction $\phi_{k,s}^{g}(u)$, i.e., $\beta_{t,k,s}^{g}=\langle\mathcal{X}_{t,s}^{g}(u)-\mu_{s}^{g}(u),\phi_{k,s}^{g}(u)\rangle$; and $e_{t,s}^{g}(u)$ denotes the model error function with a mean of zero and a finite variance, and $K_{s}^{g}<n$ is the number of retained components. The eigenfunctions represent the fixed shapes of variation (e.g., longevity shift), while the scores capture the time-varying magnitude of these shifts.
@@
$K_{s}^{g}=\operatorname*{arg\,min}_{k=1,\dots,k_{\max}}\left\{\frac{\widehat{\lambda}_{k+1,s}^{g}}{\widehat{\lambda}_{k,s}^{g}}\times\mathds{1}\Big\{\frac{\widehat{\lambda}_{k,s}^{g}}{\widehat{\lambda}_{1,s}^{g}}\geq\eta\Big\}+\mathds{1}\Big\{\frac{\widehat{\lambda}_{k,s}^{g}}{\widehat{\lambda}_{1,s}^{g}}<\eta\Big\}\right\},$ (3)
where $\mathds{1}\{\cdot\}$ is the binary indicator function, $\eta=\frac{1}{\ln[\max(\widehat{\lambda}_{1,s}^{g},n)]}$ is a small positive value. Instead of searching through $n$, we restrict our searching range by setting $k_{\max}=\#\{k|\widehat{\lambda}_{k,s}^{g}\geq\frac{1}{n}\sum^{n}_{k=1}\widehat{\lambda}_{k,s}^{g},k\geq 1\}$. For comparison, we also consider $K_{s}^{g}=6$ as suggested in Hyndman et al. (2013). Note that when $K_{s}^{g}=1$, it reduces to the Lee & Carter’s (1992) model.
@@
$\widehat{\mathcal{X}}_{n+h|n,s}^{g}(u)=\text{E}\big[\mathcal{X}_{n+h,s}^{g}(u)|\bm{\mathcal{X}}_{s}^{g}(u),\bm{B}\big]=\widehat{\mu}_{s}^{g}(u)+\sum^{K_{s}^{g}}_{k=1}\widehat{\beta}_{n+h|n,k,s}^{g}\widehat{\phi}_{k,s}^{g}(u),$
@@ 4.2 Multivariate functional time-series (MFTS) forecasting method
The univariate functional time-series method does not account for correlations between female and male data within the same region. However, explicitly modeling the correlation between multiple series may improve forecast accuracy. We consider a multivariate functional time-series method to jointly model and forecast multiple series that may be correlated.
Let $\mathcal{X}_{t,s}^{F}(u)$ and $\mathcal{X}_{t,s}^{M}(u)$ represent female and male subnational age distribution of death counts. As our multiple functional time series have the same function support, we consider data where each observation consists of $\omega\geq 2$ functions, that is, $\mathcal{X}_{t,s}^{F}(u),\mathcal{X}_{t,s}^{M}(u)]\in\mathds{R}^{2}$, where $u\in\mathcal{I}$.
Multivariate functional time series are stacked into a vector. Let $\nu^{F}_{s}(u)$ and $\nu^{M}_{s}(u)$ represent the mean function, respectively. For $u,v\in\mathcal{I}$, the theoretical cross-covariance function can be defined with elements
@@
where $\bm{\mathcal{X}}_{s}(u)=[\bm{\mathcal{X}}_{s}^{F}(u),\bm{\mathcal{X}}_{s}^{M}(u)]$, $\bm{\mathcal{X}}_{s}^{F}(u)=[\mathcal{X}_{1,s}^{F}(u),\dots,\mathcal{X}_{n,s}^{F}(u)]$ denote stacked historical functions, and $\bm{\nu}_{s}(u)=[\bm{\nu}_{s}^{F}(u),\bm{\nu}_{s}^{M}(u)]$. Moreover, $\bm{\beta}_{s}=(\bm{\beta}_{1,s},\dots,\bm{\beta}_{K,s})$ and $\bm{\beta}_{1,s}=(\beta_{1,1,s},\dots,\beta_{n,1,s})$ are the vectors of the principal component scores, and
$\bm{\Phi}_{s}(u)=\left[\begin{array}[]{cccccc}\phi_{1,s}^{F}(u)&\cdots&\phi_{K,s}^{F}(u)&&\cdots&\\ &\cdots&&\phi_{1,s}^{M}(u)&\cdots&\phi_{K,s}^{M}(u)\end{array}\right]_{2\times(K\times 2)},$
@@
Conditioning on the previous functions $\bm{\mathcal{X}}_{s}(u)$ and the estimated functional principal components $\bm{\Phi}_{s}(u)$, the $h$-step-ahead point forecast of $\bm{\mathcal{X}}_{n+h,s}(u)$ can be expressed as
@@ 4.3 Multilevel functional time-series (MLFTS) forecasting method
The multilevel functional data model strongly resembles the two-way functional analysis of variance studied by Morris & Carroll (2006) and Cuesta-Albertos & Febrero-Bande (2010). The basic idea is to extract common patterns shared by multiple series $R_{t,s}(u)$ and series-specific patterns $U_{t,s}^{g}(u)$. The common trend and series-specific trend are modeled by projecting them onto the eigenvectors of the covariance functions of the aggregated and series-specific stochastic processes, respectively. For $t=1,\dots,n$, a curve can be expressed as
@@
The sample versions of the series-specific mean function, a common trend, and the series-specific residual trend for a set of functional time series can be estimated by
@@
Plugging (5) into (7) into (4), we obtain
@@
When the common trend explains the primary mode of total variability, the within-cluster variability is close to 1.
@@
4.4 Two-way functional analysis of variance (FANOVA
To model $\mathcal{X}_{t,s}^{g}(u)$, we consider a two-way FANOVA decomposition, expressed as:
@@ 4.4 Two-way functional analysis of variance (FANOVA
The FANOVA model can be estimated by means with
@@
Some identifiability constraints exist, $\forall u\in\mathcal{I}$, $\sum^{n_{s}}_{s=1}\alpha_{s}(u)=\sum^{2}_{g=1}\beta^{g}(u)=0$, and $\sum^{n_{s}}_{s=1}\mathcal{X}_{t,s}^{g}(u)=\sum^{2}_{g=1}\mathcal{X}_{t,s}^{g}(u)=0$, $\forall t$. The FANOVA model decomposes the functional time series (FTS) jointly for all prefectures and genders into three parameters: the overall mean $\widehat{\mu}(u)$, which represents the grand effect measuring the overall baseline age distribution of deaths; $\widehat{\alpha}_{s}(u)$, which represents the functional prefecture effect measuring the time-invariant functional difference attribute to each prefecture, relative to the grand mean; and $\widehat{\beta}^{g}(u)$, which represents the functional gender effect measuring the time-invariant functional difference between genders, relative to the grand mean.
The residual functions $\widehat{\epsilon}_{t,s}^{g}(u)=\mathbf{X}_{t,s}^{g}(u)-\widehat{\mu}(u)-\widehat{\alpha}_{s}(u)-\widehat{\beta}^{g}(u)$ are time-varying, for $t=1,\dots,n$. Following the methodology of Jiménez-Varón et al. (2024), we model and forecast these residual functions using the multivariate functional time series (MFTS) approach.
@@ 4.5 Two-stage functional principal component analyses
    The first functional principal component scores from each of $P$ sets of functional time series are combined into an $P\times 1$ vector. We fit factor models to the scores to further reduce the dimension into an $r\times 1$ vector, where $r<<P$. The same is carried out for the second, third, and so on until the $p_{0}$th scores. The vector of the $P$ functional time series is reduced to an $r\times p_{0}$ matrix consisting of factors.
@@
    A univariate time series model can be fitted to each factor, and forecasts are generated. Forecast factors can be used to construct forecast functions.
@@ 5 Construction of prediction intervals
We split the data sample consisting of 49 years from 1975 to 2023 into a training set consisting of 16 years, a validation set consisting of 16 years, and a testing set consisting of 17 years. While the choice of different data splits is arbitrary, it is customary in forecasting to allocate at least one-third of the total data for evaluation (Hyndman & Athanasopoulos, 2021).
Using the data in the training sample, we implement an expanding-window forecast scheme to obtain $h$-step-ahead density forecasts in the validation set for $h=1,2,\dots,16$. For each horizon, we have a different number of curves in the validation set. For instance, when $h=1$, we have 16 years to evaluate the forecast errors, denoted by $\widehat{\epsilon}_{m,s}^{g}(u)=d_{m,s}^{g}(u)-\widehat{d}_{m,s}^{g}(u)$ for $m=1,2,\dots,M=16$; when $h=16$, we have two years to evaluate the residual functions between the samples in the validation set and their corresponding forecasts. From these residual functions, we compute the functional standard deviation, denoted by
@@
$\underline{\xi}_{\alpha}\gamma(u)\leq\widehat{\epsilon}_{m,s}^{g}(u)\leq\overline{\xi}_{\alpha}\gamma(u),$
@@
Typically, the constants $\overline{\xi}_{\alpha}$ and $\underline{\xi}_{\alpha}$ are chosen equal. By the law of large numbers, one may achieve
@@
where $M$ denotes the number of years in the validation set.
@@
In Appendix C, we also present the conformal prediction interval approach of Shafer & Vovk (2008), which does not require a tuning parameter. By considering a quantile of the absolute residual functions, the pointwise prediction interval for the out-of-sample testing set is centered on the point forecast, with its spread determined by the chosen quantile.
@@ 6.1 Expanding window scheme
An expanding window analysis of a time-series model is commonly used to assess model and parameter stability and prediction accuracy over time. The expanding window analysis assesses the stability of a model’s parameters by computing parameter estimates and their forecasts over an expanding window across the sample (for details, see Zivot & Wang, 2006, pp. 313–314).
Using the first 32 years from 1975 to 2006 in the Japanese subnational age- and sex-specific life-table death counts, we produce one- to 16-step-ahead forecasts. Through an expanding-window approach, we re-estimate the parameters in the time-series forecasting models using the first 33 years from 1975 to 2007. Forecasts from the estimated models are then produced for one- to 16-step-ahead forecasts. We iterate this process by increasing the sample size by one year until reaching the end of the data period in 2023. This process produces 16 one-step-ahead forecasts, 15 two-step-ahead forecasts, $\dots$, and one 16-step-ahead forecast. We compare these forecasts with holdout samples to assess out-of-sample forecast accuracy. In Figure 5, we display a diagram of the expanding window forecast scheme for the forecast horizon $h=1$, although we also consider other forecast horizons from $h=2$ to $16$.
[…]
@@ 6.2 Point forecast error metrics
Since life-table death counts can be considered probability density functions, we use some density-evaluation measures. These measures include the discrete version of the KLD (Kullback & Leibler, 1951), the square root of the JSD (Shannon, 1948). For two probability density functions, denoted by $d_{m+\xi}$ and $\widehat{d}_{m+\xi}$, the symmetric version of the KLD, also known as Jeffrey divergence, is defined as
$\text{KLD}(h)=\$$\text{D}_{\text{KL}}(d_{m+\xi}||\widehat{d}_{m+\xi})+\text{D}_{\text{KL}}(\widehat{d}_{m+\xi}||d_{m+\xi})$
$=\$$\frac{1}{111\times(17-h)}\sum^{16}_{\xi=h}\sum_{x=1}^{111}d_{m+\xi}(u_{x})\cdot\left[\ln d_{m+\xi}(u_{x})-\ln\widehat{d}_{m+\xi}(u_{x})\right]+$
$\frac{1}{111\times(17-h)}\sum^{16}_{\xi=h}\sum_{x=1}^{111}\widehat{d}_{m+\xi}(u_{x})\cdot\left[\ln\widehat{d}_{m+\xi}(u_{x})-\ln d_{m+\xi}(u_{x})\right],$ (9)